if you double a number and take away 20 you get 50 what is the number
Found 3 solutions by macston, Cromlix, farohw:
Answer by macston(5194)

.
x=a number
2x-20=50
2x=70
x=35
ANSWER: The number is 35.

Answer by Cromlix(4381)

Hi there,
A number = 'x'
Double the number = 2x
Take away 20 = 2x - 20
You get 50 = 2x - 20 = 50
2x - 20 = 50
Collect like terms
2x = 50 + 20
2x = 70
x = 35
Number is 35
Hope this helps :-)

Answer by farohw(175)


If you double a number and take away 20 you get 50 what is the number?

2(x) - 20 = 50
2x = 70
2x/2 = 70/2
x = 35

35 is the number you will double.